# SPF-Portrait: Towards Pure Portrait Customization with Semantic Pollution-Free Fine-tuning

### v1

@@ 3.3 Semantic-Aware Fine Control Alignment
where $D_{CLIP}$ represent the cosine similarity in CLIP embedding space. $\widehat{\mathcal{M}}$ is our final SFCM, as shown in Fig. 2 and Fig. 3 (a), it represents the precise target response regions and effectively prevents over-alignment by guiding the alignment process. Therefore, the alignment constraints in Eq. 8 and Eq. 9 can be modified as follow:
@@
$\begin{array}[]{cc}\mathcal{L}_{M\mathrm{-tex}}=\sum_{j=1}^{L}\left\|(\mathcal{F}^{j}_{ori}-\mathcal{F}^{j}_{ft})\odot(1-\widehat{\mathcal{M}})\right\|_{2},\\ \mathcal{L}_{M\mathrm{-fine}}=\sum_{j=1}^{L}\left\|(Q^{j}_{ori}-Q^{j}_{ft})\odot(1-\widehat{\mathcal{M}})\right\|_{2},\\ \end{array}$ (13)
@@
3.4 Response Enhancement via Difference Vectors
@@
For text-driven portrait customization, an effective response of the target attributes is essential for success. Therefore, to improve the model’s comprehension of the target semantics, we devise a response enhancement mechanism to bolster the performance of the target attribute. Specifically, we introduce a difference vector $\Delta$, represented by the difference between the vectors of the CLIP textual space and the CLIP visual space. By incorporating the ground truth image ${x}_{0}$, we separately calculate the difference vector $\Delta({x}_{0},E_{tar})$ between the target text and ground truth image ${x}_{0}$, as well as the difference vector $\Delta(\hat{x}_{0},E_{tar})$ between the target text and the one-step prediction $\hat{x}_{0}$, formulated as:
@@
$\begin{split}{}&\Delta(\hat{x}_{0},E_{tar})=E_{I}(\hat{x}_{0})-E_{T}(E_{tar}),\\ &\Delta(x_{0},E_{tar})=E_{I}(x_{0})-E_{T}(E_{tar}),\\ \hat{x}_{0}=&\frac{\widehat{x}_{t}}{\sqrt{\bar{\alpha}_{t}}}-\frac{\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\widehat{x}_{t},t,\tau(E_{base},E_{tar}))}{\sqrt{\bar{\alpha}_{t}}},\\ \end{split}$ (14)
@@
where the $E_{I}$ and $E_{T}$ denote the CLIP vision and text encoder, respectively, while $\hat{x}_{0}$ denotes the one-step prediction of $x_{t}$ in $t$-th timestep. Then, we constrain their similarity to enhance the response of the target semantics as:
$\begin{split}\!\mathcal{L}_{enhanced}\!=\!1\!-\!D_{CLIP}(\Delta(\hat{x}_{0},\!E_{tar}),\Delta(x_{0},\!E_{tar})).\end{split}$ (15)
@@
Figure 4: Comparison with Traditional Supervision on Image Fidelity. (a) illustrates the trend of Image-Reward (IR) and CLIP Score (CLIP-T) across fine-tuning steps. Image-Reward [52] is a metric used to evaluate image fidelity. (b) displays samples from traditional method [2] and ours.
@@
Figure 5: Qualitative Comparisons with SOTA methods. We compare ours with naive fine-tuning [43], PEFT-based methods (LoRA [19], AdaLoRA [55] ) and the decoupled methods (Tokencompose [50], Magenet [58]). Please zoom in for more details.
@@
Unlike previous work [2, 26] that directly applies cross-modal supervision in CLIP space by employing the target text to supervise the one-step prediction, formulated as:
$\begin{split}\mathcal{L}_{clip}=&1-D_{CLIP}(E_{I}(\hat{x}_{0})-E_{T}(E_{tar})),\\ \end{split}$ (16)
@@
our approach reformulates the optimization objective into difference vectors rather than image-text similarity in Eq. 16. Directly cross-modal supervision overlooks the modality representation gap, causing the model to overfit the textual description during optimization and neglecting the visual fidelity of the result image. It ultimately leads to degradation in the quality of the generated images, as illustrated in Fig. 4. In contrast, we provide an effect similar to supervision within the same modality by using the difference between cross-modal vectors, mitigating the representation discrepancy inherent in direct cross-modal supervision. It simultaneously enhances the response to target semantics while improving the fidelity and coherence of the image.
@@
$\small\!\!\!\mathcal{L}_{SPF}\!=\!\mathcal{L}_{diff}\!+\!\underbrace{\lambda_{1}\mathcal{L}_{enhacned}}_{\mathrm{\>response}}\!+\!\underbrace{\lambda_{2}\mathcal{L}_{\mathrm{M\mathrm{-text}}}\!+\!\lambda_{3}\mathcal{L}_{M\mathrm{-fine}}}_{\mathrm{alignment}},$ (17)
@@
Table 1: Quantitative Comparisons Result. Rows with a gray background indicate our ablation experiments, while rows without it represent the state-of-the-art (SOTA) methods being compared. In our specific pairwise comparison, unlike general image generation, lower FID values reflect greater consistency with the original model’s behavior. It is notable that the underlined values in “Ours (w/o SFCM)” are abnormally low, as the generated portraits exhibit over-alignment with the original portraits.
@@
Naive Fine-Tuning [43] 20.41 0.57 0.21 0.63 57.77 0.24 0.21 0.67
AdaLoRA [55] 7.38 0.40 0.39 0.80 64.86 0.23 0.24 1.10
LoRA (Best Rank) [19] 9.82 0.38 0.52 0.71 58.37 0.27 0.23 1.21
TokenCompose [50] 10.93 0.41 0.32 0.81 40.22 0.27 0.24 0.71
Magnet [58] 18.92 0.48 0.38 0.61 32.87 0.26 0.26 0.97
@@
Figure 6: Qualitative Ablation Study. We independently ablate the proposed loss and the SFCM mechanism.
Figure 7: User Study Results. The percentages indicate the proportion of users who select the corresponding method.
4 Experiments
4.1 Experimental Setup
Implementation Details. We adopt the pre-trained Stable Diffusion v1.5 model [43] with Realistic_Vision_V4.0 checkpoints. All attention maps and score maps are upsampled at a resolution of 64 $\times$ 64. The hyperparameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are set to 0.6, 0.2 and 0.1. Detailed information about experiments is provided in the Appendix.
Dataset. Our training set contains 230K diverse portrait images with novel customized attributes (e.g., skin textures, hairstyles), captioned by GPT-4o and Cambrian-1 [46]. For evaluation, we construct a test set of 5K triples, each containing: (1) an original caption, (2) its corresponding original portrait generated using Realistic_Vision_V4.0, and (3) a target caption with customized attributes.
Evaluation Metrics. In our task, we aim to evaluate three key aspects: (1) preservation of the original model’s behavior, (2) responsiveness to target semantics, and (3) overall image quality. Concretely, we employ FID [17], LPIPS [56], identity similarity (ID), CLIP Image Score (CLIP-I) [41], and segmentation consistency (Seg-Cons) to measure the consistency between the original and customized portrait. We utilize the CLIP Score (CLIP-T) [41] to evaluate responsiveness to target semantics. For comprehensive image quality assessment, we use HPSv2 [51] and MPS [57].
4.2 Qualitative Evaluation
Comparison with SOTAs. We conduct a qualitative comparison of our method with the state-of-the-art (SOTA) approaches, including PEFT-based methods such as LoRA [19] and AdaLoRA [55], decoupled text embedding methods like TokenCompose [50] and Magnet [58], as well as naive fine-tuning. We compare with them on diverse customized attributes, such as age, image style, and clothing. For each target attribute, we evaluate two cases under different random seeds. As shown in Fig. 5, although LoRA [19] and AdaLoRA [55] tend to retain original performance in some cases, their performance is extremely unstable and poor in detail alignment. For instance, in row 3, column 3, there is a noticeable change in identity, whereas in row 4, column 2, the pose of portrait has transformed completely. Magnet [58] and TokenCompose [50] naively follow the input text conditions entirely, ignoring the preservation of the original model’s behavior across all test cases. For example, in row 6&7, column 9, the customization of “pencil drawing style” results in a total alteration of the portrait. In contrast, our method purely customizes target attributes while preserving the original model’s behavior in aspects such as background, pose, and identity. It demonstrates our approach effectively mitigates semantic pollution during fine-tuning. Please refer to Appendix for more results.
4.3 Quantitative Evaluation
@@
Metric Evaluation. Tab. 1 shows the quantitative results of our methods against baselines on the test set. Our method demonstrates a substantial improvement in preservation of the original performance compared to all competitors, achieving state-of-the-art performance across all metrics. It’s notable that our method significantly outperforms competitors in “Seg-Cons”, demonstrating pixel-level alignment precision that far surpasses other approaches. The optimal CLIP-T and overall scores confirm that our method enhances the response to target semantics and achieves higher-quality portrait customization.
@@
User Study. We also conduct a user study to have a comprehensive assessment of our method. We design three dimensions for evaluation: Original Behavior Consistency (OBC), Target Attribute Responsiveness (TAR), and Aesthetic Preference (AP). We invite 32 participants from different social backgrounds, with each test session lasting about 30 minutes. Users perform pairwise comparisons between our method and competitors across three dimensions. The results are as shown in Fig. 7, our method defeat all competitors in all dimensions, especially in OBC and TAR. This highlights our ability to preserve the original model’s behavior while purely adapting to new attributes. Please refer to the Appendix for more details about user study.
4.4 Original Capability Retention
@@
To further verify that our method purely learns the customized attributes without compromising the original model and attains incremental learning, we solely utilize identical Base text to evaluate whether our method can reconstruct the original portraits after fine-tuning. As shown in Fig. 8, naive fine-tuning markedly disrupts original response patterns, while our method maintains near-identical performance to original model. For example, in the top-right case, the semantics of ‘woman’ is completely corrupted by naive fine-tuning, but we not only preserves the character but also maintains high consistency in other attributes. The outstanding reconstruction of portraits across varied scenes demonstrates our method’s substantive retention of the original model’s intrinsic capabilities.
@@
Figure 8: Reconstruction Results. The three portraits for each case are only generated by the same Base text.
4.5 Ablation Study
@@
To validate the effectiveness of different components of our method, we conduct thorough ablation studies. Qualitative results, shown in Fig. 6, indicate that the absence of $\mathcal{L}_{text-consistent}$ results in weaker alignment of Base text with the original portrait, while the lack of $\mathcal{L}_{fine-grained}$ leads to inconsistencies in detailed content, such as portrait posture. Without $\mathcal{L}_{enhanced}$, the expression of the target semantics significantly degrades that fails to follow the action of ‘holding’ and with a tendency to disrupt the spatial coherence of the ‘toy bear’, degenerating into flattened textile-like patterns. Quantitative results in ablation part of Tab. 1, further validates the conclusions drawn from the visual analysis through superior performance across all metrics. Notably, although ‘w/o SFCM’ shows superior Preservation Metrics in Tab. 1, this is due to its complete disregard for target semantics and severe over-alignment with the original portrait, shown in Fig. 6. Such outcomes represent an absolute failure in our task, which is entirely undesirable.
5 Conclusion
@@
In this paper, we propose SPF-Portrait, a novel fine-tuning framework designed to address the issue of Semantic Pollution in text-driven portrait customization. By introducing original model as a reference path and utilizing contrastive learning, we achieve the goals of purely capturing the customized semantics and enabling incremental learning. We precisely preserve the original model’s behavior and ensure an effective response to target semantics by innovatively designing a Semantic-Aware Fine Control to guide the alignment process and a response enhancement mechanism for target semantics. Extensive experiments demonstrate our method can achieve the SOTA performance. In the future, we will continue to explore adapting our framework to more broad and complex scenes, striving to achieve semantic pollution-free fine-tuning for general text-to-image and text-to-video generation.

### v3

@@ 3.3. Semantic-Aware Fine Control Alignment
where $D_{CLIP}$ represent the cosine similarity in CLIP embedding space. All attention maps are upsampled at a resolution of 64 $\times$ 64 as the same as noise map. $\widehat{\mathcal{M}}$ is our final SFCM, as shown in Fig. 3 and Fig. 4 (a), it represents the precise target response regions and effectively prevents over-alignment by guiding the alignment process.
Therefore, the alignment constraints in Eq. 6 and Eq. 7 can be modified as follow:
@@
(11) $\begin{array}[]{cc}\mathcal{L}_{M\mathrm{-tex}}=\sum_{j=1}^{L}\left\|(\mathcal{F}^{j}_{ori}-\mathcal{F}^{j}_{ft})\odot(1-\widehat{\mathcal{M}})\right\|_{2},\\ \mathcal{L}_{M\mathrm{-fine}}=\sum_{j=1}^{L}\left\|(Q^{j}_{ori}-Q^{j}_{ft})\odot(1-\widehat{\mathcal{M}})\right\|_{2},\\ \end{array}$
@@
3.4. Response Enhancement via Difference Vectors
@@
In text-to-portrait customization, an excellent response to the target semantics is essential for success. Therefore, to reinforce the model’s comprehension of the target attribute, we devise a response enhancement mechanism to improve the presentation of the target semantics. Specifically, we introduce a difference vector $\Delta$, represented by the difference between the vectors of the CLIP textual space and the CLIP visual space (Abdelfattah et al., 2023; Xue et al., 2022). By introducing the ground truth image ${x}_{0}$ (a image with target attribute), we separately calculate the difference vector $\Delta({x}_{0},E_{tar})$ between the target text and ground truth image ${x}_{0}$, as well as the difference vector $\Delta(\hat{x}_{0},E_{tar})$ between the target text and the one-step prediction $\hat{x}_{0}$, formulated as:
@@
(12) $\begin{split}{}&\Delta(\hat{x}_{0},E_{tar})=\tau_{vision}(\hat{x}_{0})-\tau_{text}(E_{tar}),\\ &\Delta(x_{0},E_{tar})=\tau_{vision}(x_{0})-\tau_{text}(E_{tar}),\\ \hat{x}_{0}=&\frac{\widehat{x}_{t}}{\sqrt{\bar{\alpha}_{t}}}-\frac{\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\widehat{x}_{t},t,\tau_{text}([E^{res}_{base},E_{tar}])}{\sqrt{\bar{\alpha}_{t}}},\\ \end{split}$
@@
where the $\tau_{vision}$ and $\tau_{text}$ denote the CLIP vision and text encoder, respectively, while $\hat{x}_{0}$ denotes the one-step prediction of $x_{t}$ in $t$-th timestep. Then, we constrain their similarity to enhance the response of the target semantics as:
(13) $\begin{split}\!\mathcal{L}_{enhanced}\!=\!1\!-\!D_{CLIP}(\Delta(\hat{x}_{0},\!E_{tar}),\Delta(x_{0},\!E_{tar})).\end{split}$
@@
Figure 5. Comparison with Traditional Supervision on Image Fidelity. (a) illustrates the trend of Image-Reward (IR) and CLIP Score (CLIP-T) across training steps. Image-Reward (Xu et al., 2023) is a metric used to evaluate image fidelity. (b) displays samples from traditional method (Avrahami et al., 2022) and ours.
@@
Figure 6. Qualitative Comparisons with SOTA methods. We compare ours with naive fine-tuning (Rombach et al., 2022), PEFT-based methods (LoRA (Hu et al., 2021), AdaLoRA (Zhang et al., 2023a) ) and the decoupled methods (Tokencompose (Wang et al., 2024b), Magenet (Zhuang et al., 2024)).
@@
Unlike previous work (Avrahami et al., 2022; Kim et al., 2022) that directly applies cross-modal supervision in CLIP space by employing the target text to supervise the one-step prediction (Yin et al., 2024), formulated as:
(14) $\begin{split}\mathcal{L}_{clip}=&1-D_{CLIP}(\tau_{vision}(\hat{x}_{0})-\tau_{text}(E_{tar})).\\ \end{split}$
@@
our approach reformulates the optimization objective into difference vectors rather than image-text similarity in Eq. 14. Directly cross-modal supervision overlooks the modality representation gap, causing the model to overfit the textual description during optimization and neglecting the visual fidelity of the result image. As illustrated in Fig. 5, it ultimately leads to degradation in the quality of the generated images. In contrast, we provide an effect similar to supervision within the same modality by using the difference between cross-modal vectors, mitigating the representation discrepancy inherent in direct cross-modal supervision. It simultaneously enhances the response to target semantics while improving the fidelity and coherence of the image.
@@
(15) $\mathcal{L}_{SPF}=\mathcal{L}_{diff}+\underbrace{\lambda_{1}\mathcal{L}_{\mathrm{M\mathrm{-text}}}+\lambda_{2}\mathcal{L}_{M\mathrm{-fine}}}_{\mathrm{alignment}}+\underbrace{\lambda_{3}\mathcal{L}_{enhacned}}_{\mathrm{\>response}},$
@@
4. Experiments
4.1. Experimental Setup
Implementation Details. We adopt the Stable Diffusion v1.5 model (Rombach et al., 2022) with Realistic_Vision_V4.0 checkpoints. The hyperparameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are set to 0.2, 0.1 and 0.6. More details about experiments are provided in the Appendix.
Dataset. Our training set contains 230K diverse portraits with new attributes (e.g., skin textures, hairstyles), captioned by GPT-4o (Achiam et al., 2023) and Cambrian-1 (Tong et al., 2024). For evaluation, we create a test set of 5K triples, each with: (1) an original caption, (2) its corresponding original portrait generated using Realistic_Vision_V4.0, and (3) a target caption of customized attributes.
Evaluation Metrics. We evaluate three key aspects: (1) preservation of the original model’s behavior, (2) responsiveness to target semantics, and (3) overall image quality. Concretely, we employ FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), identity similarity (ID), CLIP Image Score (CLIP-I) (Radford et al., 2021), and segmentation consistency (Kirillov et al., 2023) (Seg-Cons) to measure the consistency between the original and customized portraits. We use the CLIP Score (CLIP-T) (Radford et al., 2021) to evaluate responsiveness to target semantics. For overall image quality assessment, we use HPSv2 (Wu et al., 2023) and MPS (Zhang et al., 2024).
Figure 7. Qualitative Ablation Study. We independently ablate the proposed loss and the SFCM mechanism.
4.2. Qualitative Evaluation
Comparison with SOTAs. We qualitatively comparison of our method with the SOTA approaches, including PEFT-based methods such as LoRA (Hu et al., 2021) and AdaLoRA (Zhang et al., 2023a), decoupled text embedding methods like TokenCompose (Wang et al., 2024b) and Magnet (Zhuang et al., 2024), as well as naive fine-tuning. We compare with them on diverse customized attributes, such as age, image style, and clothing. For each target attribute, we evaluate two cases under different random seeds. As shown in Fig. 6 and Fig. 15, although LoRA (Hu et al., 2021) and AdaLoRA (Zhang et al., 2023a) tend to retain original behavior in some cases, their performance is extremely unstable and poor in detail alignment. For instance, in row 3, column 3, there is a noticeable change in identity, whereas in row 4, column 2, the pose of portrait has transformed completely. Magnet (Zhuang et al., 2024) and TokenCompose (Wang et al., 2024b) naively follow the input text conditions entirely, ignoring the preservation of the original model’s behavior across all test cases. For example, in row 6 & 7, column 9, the customization of “pencil drawing style” results in a total alteration of the portrait. In contrast, our method purely customizes target attributes while preserving the original model’s behavior in aspects such as background, pose, and identity. It demonstrates our approach effectively address semantic pollution during fine-tuning.
More Extensions. We provide two more extensions of our SPF-Portrait: 1) As shown in Fig 16, our method reliably performs excellently in continuous replacements and additions of target text in text-to-portrait customization. 2) In Fig 17, we demonstrate the strong potential of extending our method to the General T2I domain.
Table 1. Quantitative Comparison Results. Rows without color represent comparisons with SOTA methods, while blue rows indicate our ablation experiments. In our specific pairwise comparison, unlike general image generation, lower FID values reflect greater consistency with the original model’s behavior. Notably, the underlined values in ”Ours (w/o SFCM)” are unusually low because the generated portraits overly align with the original portraits.
@@ 4.2. Qualitative Evaluation
Naive Fine-Tuning (Rombach et al., 2022) 20.41 0.57 0.21 0.63 57.77 0.24 0.21 0.67
AdaLoRA (Zhang et al., 2023a) 7.38 0.40 0.39 0.80 64.86 0.23 0.24 1.10
LoRA (Hu et al., 2021) 9.82 0.38 0.52 0.71 58.37 0.27 0.23 1.21
TokenCompose (Wang et al., 2024b) 10.93 0.41 0.32 0.81 40.22 0.27 0.24 0.71
Magnet (Zhuang et al., 2024) 18.92 0.48 0.38 0.61 32.87 0.26 0.26 0.97
@@
4.3. Quantitative Evaluation
@@
Metric Evaluation. Tab. 1 shows the quantitative results of our methods against baselines on the test set. Our method shows substantial improvement in preserving the original behavior compared to all competitors, achieving state-of-the-art performance across all metrics. It is notable that our method significantly outperforms competitors in “Seg-Cons”, demonstrating pixel-level alignment precision that far surpasses other approaches. The optimal CLIP-T and overall scores confirm that our method enhances the response to target semantics and achieves higher-quality portrait customization.
Figure 8. User Study Results. The percentages indicate the proportion of users who select the corresponding method.
@@
User Study. We also conduct a user study to have a comprehensive assessment of our method. We design three dimensions for evaluation: Original Behavior Consistency (OBC), Target Attribute Responsiveness (TAR), and Aesthetic Preference (AP). We invite 32 participants from different social backgrounds, with each test session lasting about 30 minutes. Users perform pairwise comparisons between our method and competitors across three dimensions. The results are as shown in Fig. 8, our method defeat all competitors in all dimensions, especially in OBC and TAR. This highlights our ability to preserve the original model’s behavior while purely adapting to new attributes. Please refer to the Appendix for more details.
4.4. Analysis of the fine-tuned model
@@
To further verify that our method purely learns the customized attributes without compromising the original model and attains incremental learning, we solely use identical Base text to evaluate whether our method can reconstruct the original portraits after fine-tuning. As shown in Fig. 9, naive fine-tuning markedly disrupts original response patterns, while our method maintains near-identical performance to original model. For example, in the top-right case, the semantics of ‘woman’ is completely corrupted by naive fine-tuning, but we not only retain the character but also maintains high consistency in other attributes. The outstanding reconstruction of portraits across varied scenes demonstrates our method’s substantive retention of the original model’s intrinsic capabilities.
@@
Figure 9. Reconstruction Results. The three portraits for each case are generated by the fine-tuned model using only the same Base text.
4.5. Ablation Study
@@
To validate the effectiveness of different components of our method, we conduct thorough ablation studies. Qualitative results, shown in Fig. 7, indicate that the absence of $\mathcal{L}_{text-consistent}$ results in weaker alignment of Base text response with the original portrait, while the lack of $\mathcal{L}_{fine-grained}$ leads to inconsistencies in detailed content, such as portrait posture. Without $\mathcal{L}_{enhanced}$, the expression of the target semantics significantly degrades that fails to follow the action of ‘holding’ and with a tendency to disrupt the spatial coherence of the ‘toy bear’, degenerating into flattened textile-like patterns. Quantitative results in ablation part of Tab. 1, further validates the conclusions drawn from the visual analysis through superior performance across all metrics. Notably, although ‘w/o SFCM’ shows superior Preservation Metrics in Tab. 1, this is due to its complete disregard for target semantics and severe over-alignment with the original portrait, shown in Fig. 7. Such outcomes represent an absolute failure in our task, which is entirely undesirable.
5. Conclusion
@@
In this paper, we propose SPF-Portrait, a novel fine-tuning framework designed to address the issue of Semantic Pollution in text-to-portrait customization. By introducing original model as a reference path and utilizing contrastive learning, we achieve the goals of purely learning the customized semantics and enabling incremental learning. We precisely retain the original model’s behavior and ensure an effective response to target semantics by innovatively designing a Semantic-Aware Fine-Control Map to guide the alignment process and a response enhancement mechanism for target semantics. Extensive experiments show that our method can achieve the SOTA performance. In the future, we will continue to explore adapting our framework to more broad and complex scenes, striving to achieve semantic pollution-free fine-tuning for general text-to-image and text-to-video generation.
Appendix A Details of Our Training
Figure 10. Our overall training pipeline.
A.1. Training Stage
As shown in Fig. 10, the training process of our approach consists of two stages: fine-tuning with all the parameters updated in the first stage and contrastive learning in the second stage. In the first stage, we employ conventional fine-tuning (Rombach et al., 2022) to learn target attributes, which aims to enable the T2I model to rapidly adapt to the target concepts of our dataset. For this stage, we train the model using 8 V100 GPUs with a batch size of 8, iterating for 2 epochs and a learning rate of 1e-5. In the second stage, the training process follows the approach outlined in the main text. The goal is to enable the T2I model to grasp pure target concepts without compromising the original model’s performance, thereby preventing semantic pollution caused by the target text. Due to the additional memory consumption of the dual-branch architecture, we set the batch size to 2, iterating for 5 epochs with a learning rate of 5e-5. The same dataset and optimizer (AdamW with default parameters: beta1=0.9, beta2=0.999, weight decay=0.01) are used for both the first and second stages.
Due to the dual-path training framework in second stage, which requires an additional frozen original model compared to standard fine-tuning, our approach incurs extra memory costs and increased computation time. We provide the corresponding resource consumption details in Tab. 2. The frozen model (which doesn’t participate parameter updates) adds only 5GB of GPU memory overhead under typical FP32 precision settings.
Table 2. Computation Time and Memory Usage of Training under Different Data Type. The data in bold represents our implementation configuration.
Method Data Type
FP16 FP32 BP16
Stage-1 (w/o reference pat) 1.92s/iter (17GB) 2.28s/iter (23GB) OOM
Stage-2 (w/ reference path) 2.1s/iter (21GB) 3.26s/iter (28GB) OOM
A.2. Training Dataset
Our work focuses on preventing semantic pollution in fine-tuning portrait T2I models while enabling the model to learn the concepts from the target attributes. To achieve this, we constructed a dataset containing various image-text pairs related to portrait concepts for training the T2I diffusion model. Considering the quality and diversity of the dataset, we utilized widely adopted community checkpoints for portrait generation as the checkpoints for the Stable Diffusion (SD) model, including RealVisXL_V1.0 and HumanModel, to generate portrait images encompassing a wide range of attributes. The attribute statistics and corresponding samples are shown in Tab. 3 and Fig. 11, respectively.
To improve dataset quality, we focus on two aspects: 1) enhancing image-text alignment using FLIP (Li et al., 2024a), a CLIP checkpoint specifically for portraits, to retain the top 30% of matching pairs, and 2) improving visual fidelity by filtering images with a Human Aesthetic Preference Score (HPS) and Image-Reward (IR).
Figure 11. Example of our training datasets.
Table 3. Details of Our Training Dataset. The specific categories of character attributes covered by our training dataset.
Category number
Facial Attributes 52021
Clothing 67871
Image Style 36786
Appearance 27508
Accessories 45200
Figure 12. Comparison of results across different updated network architectures in our constraive pipeline. ”Full Weights” indicates that all network parameters are updated, ”LoRA on Cross-Atten” refers to the integration of LoRA into the Cross-Attention modules, and ”Adapter on Cross-Atten” denotes the addition of parallel cross-attention layers, akin to IP-adapter (Ye et al., 2023).
Appendix B Analysis of Fine-tuning Architecture.
Table 4. Quantitative Comparisons with other architecture.
Method Preservation Overall Responsiveness
FID ($\downarrow$) LPIPS ($\downarrow$) ID ($\uparrow$) CLIP-I ($\uparrow$) Seg-Cons ($\uparrow$) HPSv2 ($\uparrow$) MPS($\uparrow$) CLIP-T ( $\uparrow$)
Full Weights 7.82 0.40 0.309 0.81 48.39 0.22 0.87 0.26
LoRA on Cross-Atten 7.10 0.39 0.487 0.61 68.37 0.24 1.21 0.26
Adapter on Cross-Atten 5.93 0.37 0.520 0.80 61.70 0.25 1.31 0.27
Ours 4.50 0.35 0.55 0.83 75.74 0.28 1.49 0.30
During the contrastive learning of the second stage, our approach exclusively trains the parameters in the cross-attention modules. We compare results across various network architectures, including ”full-weight”, ”LoRA on cross-attention”, and ”additional adapters”. As illustrated in Fig. 12, all architectures under our contrastive learning achieve some level of alignment. Notably, ”LoRA on cross-attention”, ”Adapter on Cross-Atten” and ”Cross-Atten(ours)” outperform the ”full weights” in alignment, this is because the diffusion model relies on the cross-attention mechanism for text-conditioned control, and optimizing the most critical parameters enables a better understanding of independent target attributes. However, ”LoRA on Cross-Atten”, due to its limited learnable parameters, falls short in understanding the original behavior compared to our method. Ours achieves a superior balance between alignment and attribute learning. ”Adapter on Cross-Atten” achieves the suboptimal performance, as it independently adjusts all the parameters of cross-attention module. However, the isolated attention structure limits the interaction between target text features and base text features, rendering in partial misalignment. The results in Tab. 4 further validate our conclusions.
Appendix C Sensitivity Analysis of Loss
To determine the optimal settings for the three loss hyperparameters, we conducted a comprehensive sensitivity analysis. As shown in Fig. 13 The three segments of the plot correspond to the hyperparameters in Eq. 11 ($\lambda_{1}\rightarrow\mathcal{L}_{M-text}$, $\lambda_{2}\rightarrow\mathcal{L}_{M-fine}$, $\lambda_{3}\rightarrow\mathcal{L}_{M-enhanced}$) , demonstrating how FID scores vary with their values. Our analysis reveals that the optimal configuration occurs at $\lambda_{1}=0.2,\lambda_{2}=0.1,\lambda_{3}=0.6$, achieving the best FID score of 4.503 reported in our main results. It’s noticed that the orange dashed line indicates the FID (4.013) of ”Ours(w/o SFCM)” from Tab. 1, which exhibits over-alignment as visualized in Fig. 7.
Figure 13. Sensitivity analysis of three loss components ($\lambda_{1}\rightarrow\mathcal{L}_{M-text}$, $\lambda_{2}\rightarrow\mathcal{L}_{M-fine}$, $\lambda_{3}\rightarrow\mathcal{L}_{M-enhanced}$) with respect to FID scores.. FID varies with different parameter values for each loss component. FID=4.503 (optimal performance) and FID=4.013 (over-alignment).
Appendix D Ablation Study of Training Stage
The main contribution of our method is the addition of an extra training stage on top of naive fine-tuning. To demonstrate the effectiveness of the two-stage training strategy, we conduct an ablation study on the training stages. As shown in Tab. 5, if only the second-stage contrastive learning is used, the model struggles to learn clean target attributes, resulting in significantly poor performance on ”CLIP-T.” On the other hand, with only stage 1, the model is entirely affected by semantic pollution, failing to align with the original model behavior, thus performing worse on preservation metrics.
Table 5. Ablation Study of the training Stage.
Method Preservation Responsiveness Overall
Only Stage-1 (Naive Fine-tuning) 20.41 0.57 0.21 0.63 57.77 0.24 0.21 0.67
Only Stage-2 7.18 0.38 0.13 0.71 63.72 0.19 0.24 1.12
Stage-1&2 (Ours) 4.50 0.35 0.55 0.83 75.74 0.30 0.28 1.49
Appendix E Discussion about Editing Methods
As mentioned in the related work Sec. LABEL:sec:relat_ed_work, incorporating text-driven editing methods (Deutch et al., 2024; Wang et al., 2024c; Kim et al., 2022; Ju et al., 2024) into the T2I model pipeline can produce similar results to ours. Here, we elaborate on the distinctions between our work and editing models and demonstrate that the improvement on inversion-based editing models when replacing their T2I model with ours.
The core distinction of our work lies in preventing additional textual concepts from disrupting T2I models, which fundamentally differs from I2I editing models that primarily focus on image manipulation through precise local modifications. Although the visual results of our method are presented in a pairwise comparison which may resemble those of editing work, the purpose is to demonstrate that our incremental learning approach preserves the integrity of the original model.
For an ideal AI-driven text-to-portrait creation, users aim for text to function like a brush in traditional painting, enabling targeted modifications to specific regions while preserving others unchanged. With existing technology, users can only achieve this by combining text-driven editing models, requiring: 1) Initial creation using a T2I model, 2) Refinement with an I2I editing model. However, in our framework, the T2I model can directly modify images via controlled text input during continuous generation, eliminating the need for additional I2I editing models. It can maintain consistency across continuous generations by preserving identical content for shared text elements. This makes the creative process more controllable, convenient, and aligned with intuition.
Appendix F Details of User study
We provide more details on our user study implementation. Besides qualitative and quantitative comparisons, we also conduct a user study to determine whether our method is preferred by humans and to underst and how people perceive emotions. We invite 32 participants from different social backgrounds and each test session lasts about 30 minutes. During the investigation, as illustrated in Fig. 14, we conducted a pairwise comparison between our method and competitors across three key dimensions: Original Behavior Consistency, text alignment, and human preference. For ”Original Behavior Consistency”, users were asked to select which of the two images better preserved consistency with the original model’s outputs. For ”Target Attribute Response”, users evaluated which image more accurately reflected the target text description. For ”Aesthetic Preference”, users judged which image aligned better with their aesthetic preferences, considering factors such as visual quality and the absence of artifacts or distortions. This comprehensive evaluation framework ensures a thorough and objective assessment of our method’s performance relative to existing approaches. The generation results are evaluated on three dimensions: image fidelity, text alignment, and human preference.
Figure 14. The investigation page in user study.
Figure 15. More Results of SPF-Portrait on Text-to-Portrait Customization.
Figure 16. Results of continuous replacements and additions of target text in text-to-portrait customization. Our method demonstrates stable and excellent performance in continuous customization tasks, indicating its potential to play a role in the application scenarios of continuous AI creation.
Figure 17. Results of extending our method to the general Text-to-Image domain. These excellent experimental results demonstrate the feasibility of extending our method to the general T2I domain. Our method has the potential to address the issue of semantic pollution in fine-tuning and to achieve incremental learning within the general T2I domain.